### v2

@@ 2 Medical Applications of LLMs
Medical applications of LLMs can be defined as the intersection of two sets: tasks that LLMs can accomplish and potential healthcare needs where predominantly textual LLMs can add value. Although both sets are finite, the cardinality of the first set is arguably smaller given the vast scope of the medical domain. Thus, our approach to categorizing medical applications is based on LLMs’ tasks of various granularity, which, in turn, have a large area of intersection with NLP tasks in general. We focus on practically significant LLM tasks from an application perspective, as well as notable macro applications that encompass multiple NLP tasks, deliberately excluding specific fine-grained NLP tasks such as coreference resolution and dependency parsing. Table 3 provides an overview of selected LLM applications and their use in the medical domain. Subsequent subsections describe each task and present examples of notable implementations.
@@
Table 3: Summary of Major LLM Applications in Medical Domain
@@ 2.1 Text Generation
The task of text generation in the medical domain involves creating contextually accurate and relevant medical texts based on a sequence of prior tokens and specific context. This task may include generating clinical notes, patient reports, or drafts of research papers. The primary challenge is ensuring that the generated text is precise, medically accurate, and adheres to relevant privacy and ethical standards.
@@
To achieve this, the probabilistic framework underpinning text generation models is used. The probability of generating the next token $x_{t}$, given a sequence of previous tokens $x_{1},x_{2},\ldots,x_{t-1}$ and additional context, can be defined as:
$P(x_{t}\mid x_{1},x_{2},\ldots,x_{t-1};\text{context})$
@@
In this formulation, ‘context’ may represent various factors depending on the nature of the task. For instance, when generating patient reports, the context would include the available information about the patient, such as vital statistics, previous diagnoses and treatments.
@@
Decoder-based LLMs inherently generate text by predicting the probability of the next token based on preceding tokens and contextual information. The methods for providing context vary. The simplest approach is using prompts given directly to the LLM. However, because the knowledge within LLMs is static, integrating external knowledge sources can be advantageous. This integration is often accomplished using variations of Retrieval-Augmented Generation (RAG) techniques [172, 173].
@@
Advancements in transformer-based models have led to innovative approaches in generating clinical documentation from patient-provider interactions. A study by Brake and Schaaf compares two model designs for generating clinical notes from doctor-patient conversations using the PEGASUS-X model. The first design, GENMOD, generates the entire note in one step, while the second design, SPECMOD, generates each section independently. The study aims to evaluate the consistency of the generated notes in terms of age, gender, body part, and coherence. Evaluations were performed using ROUGE and Factuality metrics, human reviewers, and the Llama2 LLM. Results indicate that GENMOD improves consistency in age, gender, and body part references, while SPECMOD may have advantages in coherence depending on the interpretation. The study uses a proprietary dataset with 10,859 doctor-patient conversations for training and testing [174]. Nair et al. present MEDSUM-ENT, a multi-stage approach to generating medically accurate summaries from patient-provider dialogues. Using GPT-3 as the backbone, the approach first extracts medical entities and their affirmations from conversations and then constructs summaries based on these extractions through prompt chaining. The model leverages few-shot prompting and dynamic example selection to improve entity extraction and summarization. The dataset used for evaluation consists of 100 de-identified clinical encounters from a telehealth platform. MEDSUM-ENT demonstrates improved clinical accuracy and coherence in summaries compared to a zero-shot, single-prompt baseline, as evidenced by both qualitative physician assessments and quantitative metrics designed to capture medical correctness [118].
@@
The integration of LLMs with visual models facilitates the automatic generation of medical imaging reports. Chen et al. developed Dia-LLaMA [115], a framework that utilizes the LLaMA2-7B model combined with a pre-trained ViT3D [175] to manage high-dimensional CT data. It features a disease prototype memory bank and a disease-aware attention module to counteract the imbalance in disease occurrence. The framework, tested on the CTRG-Chest-548K dataset [120], showed superior performance in clinical efficacy and natural language generation metrics, surpassing other methods [115]. Another approach, R2GenGPT, enables radiology report generation by utilizing a visual alignment module that aligns visual features from chest X-ray images with the word embedding space of LLMs, thereby enhancing the capability of static LLMs to process visual data. It explores three alignment strategies—shallow, deep, and delta—each varying in trainable parameters. Evaluated on the IU-Xray and MIMIC-CXR datasets, R2GenGPT achieved impressive results in model efficiency and clinical metrics, leveraging the Swin Transformer and Llama2-7B model for enhanced integration [114].
@@
Evaluating commercial models for clinical tasks inevitably draws the interest of the scientific community. Ali et al. evaluated the use of ChatGPT for generating patient clinic letters, focusing on its readability, factual correctness, and human-like quality by testing the model with shorthand instructions simulating clinical input for creating letters addressing skin cancer scenarios. The study involved 38 hypothetical clinical scenarios, including basal cell carcinoma, squamous cell carcinoma, and malignant melanoma. Readability was assessed using the online tool Readable, targeting a sixth-grade reading level. Two independent clinicians evaluated the letters’ correctness and human-like quality using a Likert scale. The study found that ChatGPT-generated letters scored highly in correctness and human-like quality, comparable to letters written by humans, highlighting the potential of AI to enhance clinical communication while emphasizing the need for careful regulation and human oversight to mitigate risks [176].
@@ 2.2 Token Classification
Given a sequence of tokens $X=(x_{1},x_{2},\ldots,x_{n})$, the token classification task involves assigning a label $y_{i}$ from a set of categories $C$ to each token $x_{i}$ in the sequence. This can be expressed as:
$y_{i}=f(x_{i},\text{context})$
where $y_{i}$ is the label for the token $x_{i}$, and $f$ is a function mapping each token and its contextual information to a label in $C$. In this formulation, “context” usually includes the surrounding tokens and may also encompass information outside of the sequence $X$, such as external vocabulary.
Typical implementations for the token classification task involve both masked and generative language models. With masked models, such as BERT, the model is fine-tuned on a labeled dataset of clinical notes. Here, tokens are masked not for prediction but rather to infer their labels from the context, using a classification layer. In contrast, generative models typically use prompts to facilitate the generation of text already annotated with labels, thereby directly producing labeled sequences.
The widespread use of medical abbreviations and acronyms often leads to misunderstandings, necessitating accurate disambiguation of these terms to safeguard against misinterpretations that could jeopardize patient care [177]. The process of mapping the short forms of medical terms to their full expressions is referred to as clinical acronym disambiguation. Wang and Khanna evaluated the performance of various clinical BERT-based language models on the Clinical Acronym Sense Inventory (CASI) dataset and found that ClinicalBert addresses the task effectively, achieving an F1 score of 91.49% [130]. Additionally, Sivarajkumar et al. assessed the capabilities of generative large language models, including GPT3.5, BARD, and Llama2, in acronym disambiguation using the same dataset. Their study revealed that these models perform well in acronym disambiguation even without fine-tuning, with GPT3.5 achieving the highest accuracy of 0.96 [178].
While the problem of clinical acronym disambiguation might seem largely addressed, several challenges persist. For instance, Kugic et al. conducted a study on clinical acronym disambiguation in German using ChatGPT and BING, achieving an F1 score of 0.679, which highlights the need for improvement [125]. Another concern involves the datasets used in experiments. There is a possibility that LLMs may memorize specific terms, which could misrepresent their true disambiguation capabilities [179][180]. This issue necessitates further investigation to ensure the reliability of LLMs in clinical abbreviation disambiguation tasks with realistic datasets.
@@ 2.3 Sequence Classification
Sequence classification tasks in the medical domain involve assigning a categorical label to an entire sequence of text, rather than to individual tokens. This could involve classifying entire clinical documents or patient notes into categories such as diagnosis, treatment recommendation, or urgency level.
The task can be formulated as follows. Given a document $d$ that consists of a sequence of tokens $d=(x_{1},x_{2},\ldots,x_{n})$, the sequence classification function assigns a label $y$ from a set of categories $C$, expressed by the equation:
$y=f(d)$
where $y$ is the category associated with the document $d$, and $f$ is a function that maps the entire document, taking into account its contextual coherence and thematic structure, to a label in $C$. LLMs can serve as implementations of the function $f$.
Masked LLMs are particularly well-suited for sequence classification tasks due to their inherent structure. Originally, these models predict masked tokens within a sequence. For the purpose of classification, they can be adapted by employing either the output of a [CLS] token (in case of BERT-like models) or by utilizing various pooling functions on the embeddings. In the case of the [CLS] token, a fully connected layer is added on top of the token’s output, which is designed to capture the context of the entire sequence. Alternatively, pooling functions such as average pooling or max pooling can be applied to the embeddings of the sequence to aggregate information across the entire input.
Generative LLMs can be either simply prompted or fine-tuned in a way that they learn to predict class labels as part of the sequence. This can be done by training the model on inputs where a special token or delimiter indicates the end of the input and the start of the output label. Although less common, integrating a linear layer after the final token output in generative models could refine the logits corresponding to class predictions, sharpening the classification boundaries set by the model’s generative nature.
@@ 2.3.1 Suicidal Behavior Prediction
Suicidal behavior prediction tasks predominantly focus on analyzing individuals’ social media activities. Dus and Nefedov developed an automated tool for identifying potential self-harm indications in social media posts, treating the task as a binary classification problem [135]. Given an input $x$, the objective is to predict a binary label $y$, where:
$y=\begin{cases}1&\text{if suicidal behavior is detected,}\\ 0&\text{otherwise.}\end{cases}$
Here, $x$ represents features extracted from social media posts. The model estimates $P(y=1|x)$, the probability that the input suggests suicidal behavior, using a fine-tuned ELECTRA model. This model was trained on data from Kaggle’s “Suicide Watch” dataset [137] and additional social media sources. Their method attained a noteworthy accuracy rate of $93\%$ and an F1 score of $0.93$ [135].
Beyond mere social media post analysis, Levkovich et al. assessed ChatGPT-3.5 and ChatGPT-4’s ability to evaluate suicide risk based on perceived burdensomeness and thwarted belongingness. By comparing ChatGPT’s assessments to those made by mental health professionals using vignettes, they discovered that ChatGPT-4’s evaluations were in close alignment with professional judgments. In contrast, ChatGPT-3.5 tended to underestimate suicide risk, underscoring the limitations of these models in this specific area [181].
In summary, while treating suicidal behavior identification as a straightforward classification task on social media posts can lead to impressive scores using standard classification metrics, the practical and ethical implications of such approaches, including potential breaches of autonomy and principles of non-maleficence, are debatable. Well-structured vignette studies on the effectiveness of LLMs and other models can further advance research in this area. Additionally, exploring the potential of human-AI collaboration represents another promising research direction in this field.
@@ 2.3.2 Modeling Patient Timeline
Kraljevic et al. introduced Foresight, a GPT-2-based pipeline developed for modeling biomedical concepts extracted from clinical narratives. This pipeline employs NER and linking tools to transform unstructured text into structured, coded concepts. Utilizing datasets from three hospitals, covering over 800,000 patients, Foresight showed promise in forecasting future medical events. Its effectiveness was validated by clinicians on synthetic patient timelines, highlighting its potential in real-world risk forecasting and clinical research [133].
@@
Aside from LLMs, generative adversarial networks (GAN) have gained popularity, extending beyond their initial image generation domain. Shankar et al. proposed Clinical-GAN, which merges Transformer and GAN methodologies to model patient timelines, focusing on predicting future medical events based on past diagnosis, procedure, and medication codes. Tested on the MIMIC-IV dataset [139], Clinical-GAN outperformed baseline methods in trajectory forecasting and sequential disease prediction [182]. Another study employed GAN for predicting the length of stay in emergency departments. The learning process was done in multiple stages. Initially, an unsupervised training phase used a generator and discriminator to approximate the probability distribution and perform feature discovery and reconstruction. Discriminator was then fine-tuned to optimize its parameters for global optimum. A predictor layer, initially randomly initialized, was added and optimized during fine-tuning, enabling the model to map observations to their lengths of stay. The model was trained on data from the Pediatric Emergency Department in CHRU-Lille and proved the potential of GANs in this field [183].
@@
Medical outcome prediction can be seen as a subtask of modeling patient timeline and is often scoped to either predict mortality, outcomes of a specific disease, or risk of progression from one disease to another. A recent study by Shoham and Rappoport [134] examined data related to chronic kidney disease, acute and unspecified renal failure, and adult respiratory failure from the MIMIC-IV and eICU-CRD datasets. Using this data, the team generated labeled datasets for disease diagnosis prediction based on patient histories. They introduced a method named Clinical Prediction with Large Language Models (CPLLM) by fine-tuning LLMs (Llama2 and BioMedLM) using medical-specific prompts to help the models understand complex medical concept relationships. Xie et al. used EHR analysis to predict epilepsy seizures, leveraging Bio_ClinicalBERT, RoBERTa, and T5, achieving an F1 score of $0.88$ in outcome classification [146].
@@
A notable approach to predict outcomes of COVID-19 patients was proposed by Henriksson et al. [184]. The authors created a model that combines structured data and unstructured clinical notes in a multimodal fashion, leveraging a clinical KB-BERT model for multimodal fine-tuning. Trained on data from six hospitals in Stockholm, Sweden, their model effectively predicted 30-day mortality, safe discharge, and readmission of COVID-19 patients in the emergency department.

### v3

@@ 3 Medical Applications of Language Models
Medical applications of language models can be defined as the intersection of two sets: tasks that language models can accomplish and potential healthcare needs where predominantly textual language models can add value. Although both sets are finite, the cardinality of the first set is arguably smaller given the vast scope of the medical domain. Thus, our approach to categorizing medical applications is based on language models’ tasks of various granularity, which, in turn, have a large area of intersection with NLP tasks in general. We focus on practically significant tasks from an application perspective, as well as notable macro applications that encompass multiple NLP tasks, deliberately excluding specific fine-grained NLP tasks such as coreference resolution and dependency parsing.
Table 4 provides an overview of selected applications of language models in the medical domain, focusing on clinical practice and related research. Subsequent subsections delve into each task, highlighting key applications, notable datasets, and exemplary implementations. Within the scope of this study, we intentionally exclude applications in medical education and highly specialized areas of biomedical research. Readers interested in a wider range of applications of language models in medicine may refer to the paper by Thirunavukarasu et al. [10].
@@
Table 4: Summary of Major Applications of Language Models in Medical Domain
@@ 3.1 Text Generation
The task of text generation in the medical domain involves creating contextually accurate and relevant medical texts based on a sequence of prior tokens and specific context. This task may include generating clinical notes, patient reports, or treatment plans. The primary challenge is ensuring that the generated text is precise, medically accurate, and adheres to relevant privacy and ethical standards.
@@
To achieve this, text generation models rely on a probabilistic framework. The probability of generating the next token $x_{t}$, given a sequence of previous tokens $x_{1},x_{2},\ldots,x_{t-1}$ and additional context $c$, can be defined as:
$P(x_{t}\mid x_{1},x_{2},\ldots,x_{t-1};c)$
@@
In this formulation, context $c$ represents various factors depending on the nature of the task. For instance, when generating patient reports, the context would include the available information about the patient, such as vital statistics, previous diagnoses and treatments.
@@
Decoder-based language models inherently generate text by predicting the probability of the next token based on preceding tokens and contextual information. The methods for providing context vary. The simplest approach is using prompts given directly to the language model. However, because the knowledge within language models is static, integrating external knowledge sources can be advantageous. This integration is often accomplished using variations of Retrieval-Augmented Generation (RAG) techniques [194, 195].
@@
Advancements in transformer-based models have led to innovative approaches in generating clinical documentation from patient-provider interactions. A study by Brake and Schaaf [196] compares two model designs for generating clinical notes from doctor-patient conversations using the encoder-decoder-based PEGASUS-X model [197]. The first design, GENMOD, generates the entire note in one step, while the second design, SPECMOD, generates each section independently. The study aims to evaluate the consistency of the generated notes in terms of age, gender, body part, and coherence. Evaluations were performed using ROUGE[198] and Factuality[199] metrics, human reviewers, and the Llama2 LLM. Results indicate that GENMOD improves consistency in age, gender, and body part references, while SPECMOD may have advantages in coherence depending on the interpretation. The study uses a proprietary dataset with 10,859 doctor-patient conversations for training and testing [196]. Nair et al. present MEDSUM-ENT[140], a multi-stage approach to generating medically accurate summaries from patient-provider dialogues. Using GPT-3 as the backbone, the approach first extracts medical entities and their affirmations from conversations and then constructs summaries based on these extractions through prompt chaining. The model leverages ICL and dynamic example selection to improve entity extraction and summarization. The dataset used for evaluation consists of 100 de-identified clinical encounters from a telehealth platform. MEDSUM-ENT demonstrates improved clinical accuracy and coherence in summaries compared to a zero-shot, single-prompt baseline, as evidenced by both qualitative physician assessments and quantitative metrics designed to capture medical correctness [140].
@@
The integration of LLMs with visual models facilitates the automatic generation of medical imaging reports by aligning visual features from models like Swin Transformer [200] with the LLM’s word embedding space, enabling effective multi-modal understanding. Chen et al. developed Dia-LLaMA [137], a framework that utilizes the LLaMA2-7B model combined with a pre-trained ViT3D [201] to manage high-dimensional CT data. It features a disease prototype memory bank and a disease-aware attention module to counteract the imbalance in disease occurrence. The framework, tested on the CTRG-Chest-548K dataset [142], outperformed other methods in various natural language generation metrics presented in the study. Another approach, R2GenGPT [136], enables radiology report generation by utilizing a visual alignment module that aligns visual features from chest X-ray images with the word embedding space of LLMs, thereby enhancing the capability of static LLMs to process visual data. It explores three alignment strategies: shallow, deep, and delta, each varying in trainable parameters. Evaluated on the IU-Xray and MIMIC-CXR datasets, R2GenGPT achieved impressive results in model efficiency and clinical metrics, leveraging the Swin Transformer[200] and Llama2-7B model for enhanced integration.
@@
Evaluating commercial models for clinical tasks inevitably draws the interest of the scientific community. Ali et al. evaluated the use of ChatGPT for generating patient clinic letters, focusing on its readability, factual correctness, and human-like quality by testing the model with shorthand instructions simulating clinical input for creating letters addressing skin cancer scenarios. The study involved 38 hypothetical clinical scenarios, including basal cell carcinoma, squamous cell carcinoma, and malignant melanoma. Readability was assessed using the online tool Readable[202], targeting a sixth-grade reading level. Two independent clinicians evaluated the letters’ correctness and human-like quality using a Likert scale. The study found that ChatGPT-generated letters scored highly in correctness and human-like quality [203].
@@ 3.2 Token Classification
Given a sequence of tokens $X=(x_{1},x_{2},\ldots,x_{n})$ and context $c$, the task is to assign a label $y_{i}\in\mathcal{Y}$ to each token $x_{i}$, where $\mathcal{Y}$ is the set of possible categories. This can be expressed within a probabilistic framework as:
$y_{i}=\arg\max_{k\in\mathcal{Y}}P(k\mid x_{i},c),$
where $P(k\mid x_{i},c)$ is the probability of token $x_{i}$ belonging to class $k$. In this formulation, context $c$ may refer to the surrounding tokens or encompass information outside of the sequence $X$, such as external vocabulary.
Encoder-based models, such as BERT, typically excel in this category of tasks. However, they require fine-tuning on a labeled dataset to adapt to specific tasks, and fine-tuning may need to be reiterated if new information or data becomes available. In contrast, models with a decoder component, can use prompts to generate text annotated with labels, thereby directly producing labeled sequences without additional task-specific fine-tuning.
The widespread use of medical abbreviations and acronyms often leads to misunderstandings, necessitating accurate disambiguation of these terms to safeguard against misinterpretations that could jeopardize patient care [204]. The process of mapping the short forms of medical terms to their full expressions is referred to as clinical acronym disambiguation. Wang and Khanna [152] evaluated the performance of various clinical BERT-based language models on the Clinical Acronym Sense Inventory (CASI) dataset and found that ClinicalBert addresses the task effectively, achieving an F1 score of 0.915. In contrast, Sivarajkumar et al. [205] assessed the capabilities of generative LLMs, including GPT3.5, Bard (Gemini), and Llama2, in acronym disambiguation using the same dataset. Their study revealed that these models perform well in acronym disambiguation without fine-tuning, with GPT3.5 achieving the highest accuracy of 0.96.
While the problem of clinical acronym disambiguation might seem largely addressed, several challenges persist. For instance, Kugic et al.[147] conducted a study on clinical acronym disambiguation in German using ChatGPT and Bing Chat (Copilot), achieving an F1 score of 0.679, which highlights the need for improvement. Another concern involves the datasets used in experiments. There is a possibility that LLMs may memorize specific terms, which could misrepresent their true disambiguation capabilities [206][207]. This issue necessitates further investigation to ensure the reliability of LLMs in clinical abbreviation disambiguation tasks with realistic datasets.
@@ 3.3 Sequence Classification
Sequence classification tasks in the medical domain involve assigning a label or multiple labels to an entire sequence of text, rather than to individual tokens. These tasks may include classifying clinical documents or patient notes into categories such as diagnosis, treatment recommendation, or urgency level.
Most sequence classification tasks fall into the category of single-label classification, which can be formulated as follows: given a sequence of tokens $X=(x_{1},x_{2},\ldots,x_{n})$, the task is to assign a label $y\in\mathcal{Y}$, where $\mathcal{Y}$ is the set of possible categories:
$y=\arg\max_{k\in\mathcal{Y}}P(k\mid X),$
where $P(k\mid X)$ represents the probability of the sequence $X$ belonging to class $k$.
Encoder-based models are well-suited for sequence classification tasks. For example, BERT can be fine-tuned for classification tasks by adding a fully connected layer on top of the embedding of the [CLS] token, which serves as a contextual representation of the entire sequence. Alternatively, pooling functions such as average pooling or max pooling can be applied to the token embeddings of the sequence to aggregate information across the entire input. Similar to token classification tasks, sequence classification often requires re-fine-tuning when new data or updated label distributions become available.
Generative language models, which incorporate a decoder component, can be either simply prompted or fine-tuned to predict class labels as part of the sequence. Fine-tuning approach typically involves training the model on inputs where a special token or delimiter demarcates the end of the input sequence and the beginning of the output label. In some cases, integrating a linear layer after the final token output in generative models can refine the logits corresponding to class predictions, enhancing the classification boundaries produced by the model’s generative framework. However, this approach introduces additional complexity and is less commonly used.
@@ 3.3.1 Suicidal Behavior Prediction
Suicidal behavior prediction tasks predominantly focus on analyzing individuals’ social media activities. Dus and Nefedov [157] proposed an automated tool for identifying potential self-harm indications in social media posts, framing the task as a binary classification problem. The set of possible categories $\mathcal{Y}$ includes two elements: $1$, indicating the presence of suicidal behavior, and 00, indicating its absence. The input, a sequence of tokens $X$, is derived from the text of social media posts. The model estimates $P(y=1\mid X)$, the probability that the input suggests suicidal behavior, by leveraging a fine-tuned ELECTRA model. Training data included samples from Kaggle’s “Suicide Watch” dataset [159], supplemented with additional social media sources. The proposed method achieved an accuracy of $0.93$ and an F1 score of $0.93$ [157].
Beyond mere social media post analysis, Levkovich et al. [208] assessed ChatGPT-3.5 and ChatGPT-4’s ability to evaluate suicide risk based on perceived burdensomeness and thwarted belongingness. By comparing ChatGPT’s assessments to those made by mental health professionals using vignettes, they discovered that ChatGPT-4’s evaluations were in close alignment with professional judgments. In contrast, ChatGPT-3.5 tended to underestimate suicide risk, underscoring the limitations of these models in this specific area.
In summary, while treating suicidal behavior identification as a straightforward classification task on social media posts can lead to impressive scores using standard classification metrics, the practical and ethical implications of such approaches, including potential breaches of autonomy and principles of non-maleficence, are debatable [209]. Well-structured vignette studies on the effectiveness of LLMs and other models can further advance research in this area. Additionally, exploring the potential of human-AI collaboration represents another promising research direction in this field.
@@ 3.3.2 Modeling Patient Timeline
Kraljevic et al. introduced Foresight [155], a GPT-2-based pipeline developed for modeling biomedical concepts extracted from clinical narratives. This pipeline employs NER and linking tools to transform unstructured text into structured, coded concepts. Utilizing datasets from three hospitals, covering over 800,000 patients, Foresight showed promise in forecasting future medical events. Its effectiveness was manually validated by clinicians on synthetic patient timelines, highlighting its potential in real-world risk forecasting and clinical research.
@@
Among different types of models, generative adversarial networks (GANs) have gained popularity, extending their applications beyond the initial domain of image generation. Shankar et al. proposed Clinical-GAN [161], which merges Transformer and GAN methodologies to model patient timelines, focusing on predicting future medical events based on past diagnosis, procedure, and medication codes. Tested on the MIMIC-IV dataset, Clinical-GAN outperformed baseline methods in trajectory forecasting and sequential disease prediction [210]. Another study [211] employed GAN for predicting the length of stay in emergency departments. The learning process was done in multiple stages. Initially, an unsupervised training phase used a generator and discriminator to approximate the probability distribution and perform feature discovery and reconstruction. Discriminator was then fine-tuned to optimize its parameters for global optimum. A predictor layer, initially randomly initialized, was added and optimized during fine-tuning, enabling the model to map observations to their lengths of stay. The model was trained on data from the Pediatric Emergency Department in CHRU-Lille and proved the potential of GANs in this field.
@@
Medical outcome prediction can be seen as a subtask of modeling patient timeline and is often scoped to either predict mortality, outcomes of a specific disease, or risk of progression from one disease to another. A recent study by Shoham and Rappoport [156] examined data related to chronic kidney disease, acute and unspecified renal failure, and adult respiratory failure from the MIMIC-IV and eICU-CRD datasets. Using this data, the team generated labeled datasets for disease diagnosis prediction based on patient histories. They introduced a method named Clinical Prediction with Large Language Models (CPLLM) by fine-tuning LLMs (Llama2 and BioMedLM) using medical-specific prompts to help the models understand complex medical concept relationships. Xie et al. [168] used EHR analysis to predict epilepsy seizures, leveraging Bio_ClinicalBERT, RoBERTa, and T5, achieving an F1 score of $0.88$ in outcome classification.
@@
A notable approach to predict outcomes of COVID-19 patients was proposed by Henriksson et al. [212]. The authors created a model that combines structured data and unstructured clinical notes in a multimodal fashion, leveraging a clinical KB-BERT model for multimodal fine-tuning. Trained on data from six hospitals in Stockholm, Sweden, their model effectively predicted 30-day mortality, safe discharge, and readmission of COVID-19 patients in the emergency department, as measured by AUC .